OpenStudy (o.lee14):

Which represents the solution(s) of the graphed system of equations, y = x2 + x – 2 and y = –2x + 2? (–2, 0) and (0, 1) (0, –2) and (1, 0) (–2, 0) and (1, 0) (0, –2) and (0, 1)

OpenStudy (quantom):

I GOT U

OpenStudy (o.lee14):

YAY

OpenStudy (quantom):

k just one question

OpenStudy (quantom):

what property are you going to use to solve this?

OpenStudy (quantom):

graphing or subsition?

OpenStudy (o.lee14):

graphing i think

OpenStudy (quantom):

k no problem one second im solving this

OpenStudy (o.lee14):

tysm!

OpenStudy (quantom):

wait one second can you write the question again?

OpenStudy (quantom):

just the question not the answers

OpenStudy (o.lee14):

Which represents the solution(s) of the graphed system of equations, y = x2 + x – 2 and y = –2x + 2?

jhonyy9 (jhonyy9):

can you solve these two equations ?

OpenStudy (quantom):

i can

jhonyy9 (jhonyy9):

not you sorry the asker i asked ?

jhonyy9 (jhonyy9):

lee do you can solve these ?

OpenStudy (o.lee14):

solve what?

jhonyy9 (jhonyy9):

x^2 +x -2 = 0 you can factoriz it cpmpletly

OpenStudy (o.lee14):

sure

jhonyy9 (jhonyy9):

do you know the way how ?

OpenStudy (quantom):

ok so its definitely not A or D

OpenStudy (o.lee14):

the factored form is x = 1, -2

OpenStudy (quantom):

I believe it is B

OpenStudy (o.lee14):

okay thank you!

jhonyy9 (jhonyy9):

the factored form of first is y=(x+2)(x-1) yes ?

OpenStudy (o.lee14):

yup

OpenStudy (o.lee14):

you think you can do one more? One of the solutions of the system of equations shown in the graph has an x-value of –4. What is its corresponding integer y-value?

OpenStudy (quantom):

ok sure :)

jhonyy9 (jhonyy9):

-4 is the x value of what equatuion ?

OpenStudy (quantom):

Can you paste the graph?

OpenStudy (o.lee14):

sure

OpenStudy (quantom):

-3

OpenStudy (quantom):

wait i have a question...

OpenStudy (quantom):

you know your first question... Was one of its answers (1,0) (-4, 10)?

OpenStudy (o.lee14):

yes? i just submitted the test though ahh i didn't pass, but thank you

OpenStudy (o.lee14):

nope

OpenStudy (quantom):

Did u get it right?

OpenStudy (quantom):

If not then it was C

OpenStudy (quantom):

So did you get them right?

OpenStudy (o.lee14):

it was right